OpenStudy (anonymous):

Find all polar coordinates of point P where P = (9,3/2 π).

OpenStudy (anonymous):

I don't quite get the question... but an angle is equal when you add a multiple of \(2\pi\) to it....\[ P_n=\left(9,\frac32\pi+2\pi n\right),\,n\in \mathbb{Z} \]

OpenStudy (anonymous):

(9, π/5 + 2nπ) or (-9, π/5 + 2nπ) (9, π/5 + (2n + 1)π) or (-9, π/5 + 2nπ) (9,π/5 + nπ) or (-9, π/5 + nπ) (9, π/5 + 2nπ) or (-9, π/5 + (2n + 1)π)

OpenStudy (anonymous):

those are the answer choices

OpenStudy (anonymous):

Okay, so if you make the magnitude of a polar coordinate negative, then it basically just shifts the magnitude over by \(\pi\).

OpenStudy (anonymous):

?? lol im sorry i dont understand

OpenStudy (anonymous):

oops, I mean it rotates the angle by \(\pi\)

OpenStudy (anonymous):

5/2

OpenStudy (anonymous):



OpenStudy (anonymous):